SWITCHlan Trouble Tickets

Transparency is very important to us. We therefore publish all trouble tickets about issues that have an impact on our customers. We show both currently open tickets as well as all closed tickets.

Ticket 20171022_1

Ticket Number: 20171022_1Ticket State: CLOSED
Ticket Opened: 2017-10-22 12:05Ticket Closed: 2017-10-23 17:24
Ticket Description: optical node at EPFL down

Problem Description:

All optical links of oEL are down since 06:27.
All ECI chassis are not reachable anymore.
No services are currently affected - all traffic is rerouted to backup path.

Actions:


2017-10-23 15:30
Both defective modules have been replaced.

2017-10-23 10:30
Two DC converter modules will have to be replaced; the replacement is scheduled for mid afternoon.

2017-10-23 08:20
All links are up, normal routing restored. We are debugging remaining problems on the rectifiers

2017-10-23 07:20
some links to neighbouring sites came up after power to the rack has been restored; we re working on the remaining problems

2017-10-22 13:31
We guess there is a power problem in Rack M01 where our optical equipment is installed.
We will ask our contacts at EPFL to check it Monday morning.


For all questions about this ticket, please send mail to noc@switch.ch
or call +41 44 268 15 30.